If nominal GDP rose from $6,500 billion in the base year to $7,030.4 billion in the current year and the current year GDP Deflator is 104, then the economy's growth rate is:

a. approximately 2.3%

b. 4%

c. 6.4%

d. approximately 10.6%
Answer #1

Nominal GDP and real GDP are same in the base year.

In real GDP in the current year is

(Nominal GDP/Real GDP)x100= Deflator

Nominal GDP in the current year is $7030.4

Deflator=104

Real GDP=$7030.4/104=$6760

Rate of growth=(( $6760-$6500)/6500) x100= (260/6500) x100= 4%

Answer is b.